# Segre forms of singular metrics on vector bundles and Lelong numbers

### v1

@@ 6. Segre forms and endomorphisms of qas bundles
Assume that $E\to X$ is a holomorphic vector bundle equipped with a locally bounded qas metric $h$. The Segre endomorphism induced by $h$ on ${\mathcal{QZ}}(X)$ is independent of the choice of reference metric, and may thus be denoted by ${\bf s}(E,h)$.
Locally, there exists a sequence of smooth metrics $h_{\epsilon}$ decreasing to $h$ when $\epsilon\to 0$ such that
$s(E,h_{\epsilon})\wedge\mu\to{\bf s}(E,h)\wedge\mu,\quad{\epsilon}\to 0,\quad\mu\in{\mathcal{QZ}}(X).$
If $h$ is of the form $h_{1}\oplus h_{2}$, then $h_{\epsilon}=(h_{1})_{\epsilon}\oplus(h_{2})_{\epsilon}$, where $(\ )_{\epsilon}$ denotes such a regularization of the corresponding metric.
In case $h$ is Griffiths negative, then such a regularization may be obtained from [9, Proposition 3.1] by representing $h$ as a matrix in a given trivialization of $E$, and letting $h_{\epsilon}=h\ast\rho_{\epsilon}$, where $\ast$ denotes the convolution with respect to a choice of local coordinates $x$, and $\rho_{\epsilon}(x)$ is a family of radial mollifiers, and it will follow by the Bedford-Taylor theory and the definition of ${\bf s}(E,h)$ that the corresponding currents converge. If $h$ is just qas, then one may instead take
(6.7) $h_{\epsilon}=h\circ\rho_{\epsilon}+A{\epsilon}I,$
where $A$ is chosen to be a large enough constant to make $h_{\epsilon}$ decreasing in ${\epsilon}$ when ${\epsilon}\to 0$, and $I$ denotes the identity matrix.
@@
As remarked above, $s^{\prime}(E,h,h_{0})$ is independent of the choice of reference metric $h_{0}$. If $Z=\emptyset$, then $s(E,h,h_{0})=s^{\prime}(E,h,h_{0})$, and hence $s(E,h,h_{0})$ is also independent of $h_{0}$.
Fix a trivialization of $E$, and a local coordinate system $x$ on $X$ such that locally, $E\cong{\mathbb{C}}^{r}_{\alpha}\times{\mathbb{C}}^{n}_{x}$. Define $h_{\epsilon}$ by (6.7). Note that $h_{\epsilon}$ depends on both the trivialization of $E$ and the local coordinate system on $X$. It is readily verified that a convolution of positive definite matrices is positive definite, and the entries are smooth, so $h_{\epsilon}$ indeed defines locally a smooth metric on $E$.
To prove the convergence result, assume that we are on the subset where $\alpha_{j}\neq 0$. With respect to the frame induced by $\alpha/\alpha_{j}$ of $L_{E}$, the metric $e^{\psi_{\epsilon}}$ induced by $h_{\epsilon}$ is represented by the function $\psi_{\epsilon}=\log|\alpha/\alpha_{j}|_{h_{\epsilon}(x)}$. Then $\psi_{\epsilon}$ is exactly of the form (2.6) with $\psi=\log|\alpha/\alpha_{j}|^{2}_{h(x)}$, and $v=\log(|\alpha/\alpha_{j}|^{2})$. By Proposition 2.3, if $A$ is chosen large enough, $\psi_{\epsilon}$ may locally be decomposed as $u_{\epsilon}+g$, where $u_{\epsilon}$ and $g$ are smooth and $u_{\epsilon}$ decrease pointwise to a locally bounded function $u$. It then follows by the Bedford-Taylor theory that
$s(L_{E},e^{\psi_{\epsilon}})\wedge\pi^{*}\mu=\sum_{k}(dd^{c}\psi_{\epsilon})^{k+r-1}\wedge p^{*}\mu\to\sum_{k}(dd^{c}\psi)^{k+r-1}\wedge p^{*}\mu.$
Taking the pushforward to $X$ finishes the proof of the convergence.
The statement about the case when $h=h_{1}\oplus h_{2}$ follows readily by noting that a convolution preserves a direct sum. ∎
Proof of Proposition 1.3.
Proposition 1.3 follows immediately from Proposition 6.4 by taking $\mu=1$. ∎
Here is our main theorem on the Segre endomorphism ${\bf s}(E)$.
@@
Assume that $E\to X$ is a holomorphic vector bundle equipped with a qas metric $h$ and a smooth reference metric $h_{0}$, and let $Z$ be the unbounded locus of $h$. Let ${\bf s}(E,h,h_{0})$ be the associated Segre endomorphism on ${\mathcal{QZ}}(X)$.
@@
(iii) If $\tau\colon X^{\prime}\to X$ is a proper mapping, then $\tau^{*}h$ is qas and,
@@
(6.11) ${\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0})\wedge\mu=s(E^{\prime},h^{\prime})\wedge{\bf s}(E,h,h_{0})\wedge\mu,\quad\mu\in{\mathcal{QZ}}(X).$
@@
(vii) Assume that $E_{1}$ and $E_{2}$ are qas bundles, and let $Z_{E_{j}}$ denotes the unbounded locus of $Z_{E_{j}}$. For any $k,\ell$, there is a normal current $w$, which is the push-forward of a current on $Z_{E_{1}}\cap Z_{E_{2}}$ such that
$({\bf s}_{k}(E_{1})\wedge{\bf s}_{\ell}(E_{2})-{\bf s}_{\ell}(E_{2})\wedge{\bf s}_{k}(E_{1}))\wedge\mu=dd^{c}w.$
@@
(i) If $h_{0}$ and $h_{0}^{\prime}$ are two smooth reference metrics, then by the first part of (vi), $({\bf s}(E,h,h_{0})\wedge\mu-{\bf s}(E,h,h_{0}^{\prime}))\wedge\mu$ belongs to ${\mathcal{QZ}}_{0}(X)$. Since ${\bf s}(E,h,h_{0})$ and ${\bf s}(E,h,h_{0}^{\prime})$ coincide outside of $Z$, it follows by Remark 4.15 that there is a current $w$, which is the pushforward of a current on $Z$ such that
(6.12) $\big{(}{\bf s}_{k}(E,h,h_{0})-{\bf s}_{k}(E,h,h_{0}^{\prime}))\wedge\mu=dd^{c}w.$
In particular, if $\mu$ has bidegree $(q,q)$, then $w$ has bidegree $(k+q-1,k+q-1)$, and hence it vanishes if $k+q\leq{\rm codim\,}Z$.
(ii) Note that if $\mu$ has bidegree $(q,q)$, then the current $w$ in (vii) has bidegree $(k+\ell+q-1,k+\ell+q-1)$. Since $w$ is the push-forward of a current on $Z$, if $k+\ell+q\leq{\rm codim\,}Z$, then $w$ vanishes. ∎
@@
Now (ii) follows since it holds when the metric is smooth, and hence also when it is locally bounded qas by Proposition 6.4.
Notice that a holomorphic map $\tau\colon X^{\prime}\to X$ induces a holomorphic map $\tau\colon{\mathbb{P}}(\tau^{*}E)\to{\mathbb{P}}(E)$. Part (iii) now follows from the definition (6.2) and Proposition 5.3 (iii).
@@
If the metric on $E$ is identically 00, i.e., $Z^{\prime}={\mathbb{P}}(E)$, then ${\bf s}(E,h,h_{0})={\bf s}(E,h_{0})$ and then the statement is trivial. Let us thus assume that $Z^{\prime}$ is not equal to ${\mathbb{P}}(E)$. It may very well happen anyhow that $Z=X$. Notice that $p^{*}\gamma$ is a product as in (4.1) on ${\mathbb{P}}(E)$. Let $\psi$ and $\psi_{0}$ be as above. Then $\psi-\psi_{0}$ is a global function on ${\mathbb{P}}(E)$ and, cf. Proposition 4.8,
@@
To prove (v), consider the metric $h_{\epsilon}^{\prime\prime}:=(h^{\prime}\oplus h)+{\epsilon}(h^{\prime}\oplus h_{0})$, which has empty unbounded locus. By Proposition 6.4, we may locally find smooth metrics $h_{{\epsilon},{\epsilon}^{\prime}}$ such that ${\bf s}(E,h_{{\epsilon},{\epsilon}^{\prime}})\to{\bf s}(E,h_{\epsilon}^{\prime\prime})$ when ${\epsilon}^{\prime}\to 0$ and $h_{{\epsilon},{\epsilon}^{\prime}}=((1+{\epsilon})h^{\prime})_{{\epsilon}^{\prime}}\oplus(h+{\epsilon}h_{0})_{{\epsilon}^{\prime}}$, where $(\ )_{{\epsilon}^{\prime}}$ denotes the regularization given by Proposition 6.4 for the corresponding metric. For each ${\epsilon}^{\prime}>0$ we have
$s(E,h_{{\epsilon},{\epsilon}^{\prime}})=s(E,((1+{\epsilon})h^{\prime})_{{\epsilon}^{\prime}})\wedge s(E,(h_{\epsilon})_{{\epsilon}^{\prime}}).$
Multiplying this equality of smooth forms with $\mu$ and letting ${\epsilon}^{\prime}\to 0$, it follows that
(6.16) ${\bf s}(E,h_{\epsilon})\wedge\mu=s(E,(1+{\epsilon})h^{\prime})\wedge s(E,h_{\epsilon})\wedge\mu.$
@@
By (i) the left hand side tends to ${\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0})\wedge\mu$. Notice that $s(E,(1+{\epsilon})h^{\prime})=s(E,h^{\prime})$. The right hand side of (6.16) is therefore equal to ${s(E,h^{\prime})\wedge s(E,h_{\epsilon})\wedge\mu}$ that tends to $s(E,h^{\prime})\wedge{\bf s}(E,h,h_{0})\wedge\mu$ by (i).

### v2

@@ 6.1. The case when $h$ is qas and non-degenerate
Let $\psi(x,y)$ be a bounded qpsh function on $U\subset{\mathbb{C}}^{n}_{x}\times{\mathbb{C}}^{m}_{y}$, such that $\psi$ is smooth in $y$ for each fixed $x$, let $v(x,y)$ be a smooth psh function, and let $\rho_{\epsilon}(x)=\rho(x/{\epsilon})/{\epsilon}^{n}$ be a family of radial mollifiers with respect to the variable $x$. Define
where $\psi*_{x}\rho_{\epsilon}$ denotes the convolution with respect to $x$. Then for any $V\subset\subset U$, $\psi_{\epsilon}$ is smooth on $V$ for small ${\epsilon}>0$, and if $A\gg 1$, then $\psi_{\epsilon}$ decreases to $\psi$ when ${\epsilon}\to 0$. Moreover,
(6.10) $(dd^{c}\psi_{\epsilon})^{k}\wedge\mu\to(dd^{c}\psi)^{k}\wedge\mu,\quad\mu\in{\mathcal{QZ}}(U).$
Since $\psi$ is locally bounded we write $(dd^{c}\psi)^{k}\wedge\mu$ rather than $dd^{c}\psi]^{k}\wedge\mu$.
@@
We can choose a constant $B>0$ such that $\psi+g$ is psh on $V$ if $g=B(|x|^{2}+|y|^{2})$. For ${\epsilon}>0$,
$\psi_{\epsilon}=\log\big((e^{\psi}*_{x}\rho_{\epsilon})e^{g}+e^{v+\log({\epsilon}A)+g}\big)-g.$
We claim that the first term $\psi^{\prime}$ on the right is psh. To see this first assume that $\psi$ is smooth. If we approximate the convolution with a Riemann sum we get
(6.11) $\log\big(\sum_{k}e^{\psi(\cdot-x_{k},\cdot)+c_{k}+g}+e^{v+\log({\epsilon}A)+g}\big).$
Since $dd^{c}g$ is translation invariant each $\psi(\cdot-x_{k},\cdot)+g$ is psh. Also $v+g$ is psh so the function in (6.11) is psh. Since such functions tend uniformly to $\psi^{\prime}$ we conclude that $\psi^{\prime}$ is psh. If $\psi$ is not smooth we approximate $u=\psi+g$ by a decreasing sequence $u_{j}$ of smooth psh functions and let $\psi_{j}=u_{j}-g$. By the same argument, each $\psi^{\prime}_{j}$ is psh, and since $\psi_{j}$ decreases to $\psi$, $\psi^{\prime}$ is psh. Thus the claim is proved.
Note that $\psi_{\epsilon}$ is smooth as a function of $x$ and $y$. Clearly $\psi_{\epsilon}\to\psi$ when ${\epsilon}\to 0$. We now prove that $\psi_{\epsilon}$ decreases with $\epsilon$} on $V$ if $A$ is large enough. To see this, we may fix $y$ and consider everything as a function of $x$, as long as the choice of $A$ is uniform in $y$. We write $\psi=u+g$, where $u$ is psh and $g$ is smooth on $X$. Then
$e^{\psi_{\epsilon}}(x)=e^{g(x)}(e^{u}\ast\rho_{\epsilon})(x)+\int(e^{g(x-x^{\prime})}-e^{g(x)})\rho_{\epsilon}(x^{\prime})dV(x^{\prime})+A\epsilon e^{v}.$
Since $e^{u}$ is psh, the first term of the right hand side decreases when $\epsilon$ decreases. To see that the whole expression is decreasing when $\epsilon$ decreases it thus suffices to prove that the middle term is differentiable in $\epsilon$, with a derivative less than $Ae^{v}$. In fact, $(d/d\epsilon)\rho_{\epsilon}(x^{\prime})={\mathcal{O}}(1/\epsilon^{n+1})$, and has support where $|x^{\prime}|\leq\epsilon$. Since $|e^{g(x-x^{\prime})}-e^{g(x)}|\leq C\epsilon$ for $|x^{\prime}|\leq\epsilon$ (with $C$ depending on the $C^{1}$-norm of $e^{g}$), the derivative of the middle term is bounded by a constant that is smaller than $Ae^{v}$ on $V$ if $A$ is chosen large enough. This choice can also be made uniformly in $y$.
It remains to prove (6.10). As before we have that $\psi_{\epsilon}=\psi_{\epsilon}+g-g=u_{\epsilon}+g$, where $u_{\epsilon}$ are psh and decrease to $u$ when ${\epsilon}\to 0$, and $g$ is smooth. Applying the Bedford-Taylor theory of convergence of Monge-Ampère products for bounded psh functions, one obtains (6.10) in case $\mu$ has the form (4.1). If $\mu=\tau_{*}\gamma$, where $\tau\colon W\to U$ is proper, then $\tau^{*}\psi_{\epsilon}+\tau^{*}g$ is psh and smooth on $\tau^{-1}V$ and decreases to $\tau^{*}\psi+\tau^{*}g$ when ${\epsilon}\to 0$. We conclude that $(dd^{c}\tau^{*}\psi_{\epsilon})^{k}\wedge\gamma\to(dd^{c}\tau^{*}\psi)^{k}\wedge\gamma$ and taking $\tau_{*}$ we get (6.10). ∎
We will now see that, as long as qas bundles have empty degeneracy loci, their Segre forms in many ways behave as in the smooth case.
@@
Assume that $E\to X$ is a holomorphic vector bundle equipped with a non-degenerate qas metric $h$.
(i) The Segre endomorphism on ${\mathcal{QZ}}(X)$ induced by $h$ is independent of the choice of reference metric, and is denoted by ${\bf s}(E,h)$. On the (possibly empty) open subset where $h$ is smooth, ${\bf s}(E,h)$ coincides with multiplication by the Segre form of $h$ there.
(ii) Locally there exist smooth metrics $h_{\epsilon}$, decreasing to $h$ when $\epsilon\to 0$, such that
(6.12) $s(E,h_{\epsilon})\wedge\mu\to{\bf s}(E,h)\wedge\mu,\quad{\epsilon}\to 0,\quad\mu\in{\mathcal{QZ}}(X).$
(iii) If $(E^{\prime},h^{\prime})$ is another qas bundle and $h^{\prime}$ is non-degenerate as well, then
(6.13) ${\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h)\wedge\mu={\bf s}(E,h)\wedge{\bf s}(E^{\prime},h^{\prime})\wedge\mu,\quad\mu\in{\mathcal{QZ}}(X).$
(iv) For any $\ell$ and $k$ the endomorphisms ${\bf s}_{k}(E,h)$ and ${\bf s}_{\ell}(E^{\prime},h^{\prime})$ commute.
Proof.
The hypothesis means that $Z=\emptyset$ and thus $s(E,h,h_{0})=s^{\prime}(E,h)$. Therefore it is independent of $h_{0}$, cf. (6.7). It follows by (2.3), (4.17) and (6.7) that if $h$ is smooth on an open set, then ${\bf s}(E,h)$ coincides with multiplication by the Segre form of $h$ there.
We now prove (ii). Fix a trivialization of $E$, and a local coordinate system $x$ on $X$ such that locally, $E\cong{\mathbb{C}}^{r}_{\alpha}\times{\mathbb{C}}^{n}_{x}$. Let $\rho_{\epsilon}$ be a family of radial mollifiers, and let
(6.14) $h_{\epsilon}=h*\rho_{\epsilon}+A{\epsilon}I,$
where $I$ is the identity matrix and $A$ is a large constant. Note that $h_{\epsilon}$ depends on both the trivialization of $E$ and the local coordinate system on $X$. It is readily verified that a convolution of positive definite matrices is positive definite, and the entries are smooth, so that each $h_{\epsilon}$ indeed defines a smooth metric on $E$.
Fix a point $x_{0}\in X$ and a small neighborhood $V\subset X$. Consider the subset of $p^{-1}V\subset{\mathbb{P}}(E)$ where $\alpha_{j}\neq 0$. With respect to the frame $\alpha/\alpha_{j}$ of $L_{E}$, the metric $e^{\psi_{\epsilon}}$ induced by $h_{\epsilon}$ is represented by the function $\psi_{\epsilon}=\log|\alpha/\alpha_{j}|_{h_{\epsilon}(x)}$. Then $\psi_{\epsilon}$ is of the form (6.9) with $\psi=\log|\alpha/\alpha_{j}|^{2}_{h(x)}$, and $v=\log(|\alpha/\alpha_{j}|^{2})$. By Proposition 6.4, if $A$ is chosen large enough, $\psi_{\epsilon}$ is decreasing when ${\epsilon}\to 0$. By compactness the same $A$ works where $\alpha_{j}\neq 0$ for any $j$. We conclude that $h_{\epsilon}$ decreases to $h$. It follows from Proposition 6.4 that
(6.15) $s(L_{E},e^{\psi_{\epsilon}})\wedge p^{*}\mu=\sum_{k}(dd^{c}\psi_{\epsilon})^{k+r-1}\wedge p^{*}\mu\to\sum_{k}(dd^{c}\psi)^{k+r-1}\wedge p^{*}\mu.$
Applying $p_{*}$ we get (6.12), keeping in mind, cf. (4.17), that ${\bf s}(E,h_{\epsilon})\wedge\mu=s(E,h_{\epsilon})\wedge\mu$ since $h_{\epsilon}$ are smooth. Thus (ii) is proved.
We now consider (iii). Let $\psi^{\prime}=\log|\alpha^{\prime}|^{2}_{h^{\prime}(x)}$. We can assume that $g$ is such that also $\psi^{\prime}+g$ is psh. Let $h^{\prime}_{{\epsilon}^{\prime}}$ be a similar regularization. It then follows that $\log(e^{\psi^{\prime}_{{\epsilon}^{\prime}}}+e^{\psi_{\epsilon}})+g$ is psh for each $\epsilon$$^{\prime}$ and $\epsilon$} and decreases to $\log(|\alpha^{\prime}|^{2}_{h^{\prime}(x)}+|\alpha|^{2}_{h(x)})+g$ when $({\epsilon}^{\prime},{\epsilon})$ decreases to $(0,0)$. With the same argument as in the proof of (ii) we conclude that
(6.16) $s(E^{\prime}\oplus E,h^{\prime}_{{\epsilon}^{\prime}}\oplus h_{\epsilon})\wedge\mu\to{\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h)\wedge\mu$
unconditionally when ${\epsilon}^{\prime},{\epsilon}\to 0$. With the product formula for smooth metrics
$s(E^{\prime}\oplus E,h^{\prime}_{{\epsilon}^{\prime}}\oplus h_{\epsilon})\wedge\mu=s(E,h_{\epsilon})\wedge s(E^{\prime},h^{\prime}_{{\epsilon}^{\prime}})\wedge\mu.$
If we first let ${\epsilon}^{\prime}\to 0$ for fixed $\epsilon$} and then let ${\epsilon}\to 0$ we get (6.13) from (6.16) and (6.12).
Part (iv) requires an additional argument; it is indeed an immediate consequence of Theorem 6.7 (vii) below. ∎
Let $E\oplus\cdots\oplus E$ be the direct sum of $k$ copies of $E$. By repeated use of (iii) we have ${\bf s}(E\oplus\cdots\oplus E,h\oplus\cdots\oplus h)={\bf s}(E,h)^{k}$ and clearly ${\bf s}(E\oplus\cdots\oplus E,h_{\epsilon}\oplus\cdots\oplus h_{\epsilon})={\bf s}(E,h_{\epsilon})^{k}$. Applying (ii) to the direct sum, noting that $(h\oplus\cdots\oplus h)_{\epsilon}=h_{\epsilon}\oplus\cdots\oplus h_{\epsilon}$, we conclude that
(6.17) ${\bf s}(E,h_{\epsilon})^{k}\to{\bf s}(E,h)^{k}.$
Since $s(E,h)={\bf s}^{\prime}(E,h)1$ it follows, cf. (6.7) that
(6.18) $s(E,h)=p_{*}\sum_{\ell=0}^{\infty}(dd^{c}\psi)^{\ell},$
which is precisely the definition (2.3) when $h$ is smooth but with the Monge-Ampère products interpreted in the Bedford-Taylor sense. Letting $\mu=1$, (ii) implies that
(6.19) $s(E,h_{\epsilon})\to s(E,h).$
Part (iv) means that $s(E\oplus E^{\prime},h\oplus h^{\prime})={\bf s}(E^{\prime},h^{\prime})\wedge s(E,h).$
Remark $6.$
If $h$ is Griffiths negative, then one can take $A=0$ in (6.14). The resulting approximation, in particular (6.19), appeared in [17]. ∎
6.2. The case with a general qas metric $h$
Here is our main theorem on the Segre endomorphism ${\bf s}(E,h,h_{0})$.
Theorem 6.7.
Assume that $E\to X$ is a holomorphic vector bundle equipped with a qas metric $h$ and a smooth reference metric $h_{0}$, let $Z$ be the degeneracy locus of $h$, and let ${\bf s}(E,h,h_{0})$ be the associated Segre endomorphism on ${\mathcal{QZ}}(X)$.
@@ 6.2. The case with a general qas metric $h$
(iii) If $\tau\colon X^{\prime}\to X$ is a proper mapping, then $\tau^{*}h$ is qas and
@@
(v) If $E^{\prime}$ is another holomorphic vector bundle, equipped with a smooth metric $h^{\prime}$, then
(6.23) ${\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h,h^{\prime}\oplus h_{0})\wedge\mu={\bf s}(E,h,h_{0})\wedge{\bf s}(E^{\prime},h^{\prime})\wedge\mu,\quad\mu\in{\mathcal{QZ}}(X).$
@@
(vii) Assume that $(E_{1},h_{1})$ and $(E_{2},h_{2})$ are qas bundles, with degeneracy loci $Z_{1}$ and $Z_{2}$, respectively, and $h_{10}$ and $h_{20}$ be smooth reference metrics on these bundles. Let $X^{\prime}\subset\subset X$ be an open relatively compact subset. For any $k,\ell$ and $\mu\in{\mathcal{GZ}}(Y^{\prime})$, there is a current $w$ on $X^{\prime}$, which is the pushforward of a current on $Z_{1}\cap Z_{2}\cap X^{\prime}$, such that
$({\bf s}_{k}(E_{1},h_{1},h_{10})\wedge{\bf s}_{\ell}(E_{2},h_{2},h_{20})-{\bf s}_{\ell}(E_{2},h_{2},h_{20})\wedge{\bf s}_{k}(E_{1},h_{1},h_{10}))\wedge\mu=dd^{c}w.$
@@
(i) If $h_{0}$ and $h_{0}^{\prime}$ are two smooth reference metrics, then by the first part of (vi),
(6.24) $\big({\bf s}_{k}(E,h,h_{0})-{\bf s}_{k}(E,h,h_{0}^{\prime}))\wedge\mu,\quad\mu\in{\mathcal{QZ}}(X),$
belongs to ${\mathcal{QZ}}_{0}(X)$. Since ${\bf s}(E,h,h_{0})$ and ${\bf s}(E,h,h_{0}^{\prime})$ coincide outside of $Z$, it follows from Proposition 4.15 that if $\mu$ has bidegree $(q,q)$, then (6.24) vanishes if $k+q-1<{\rm codim\,}Z$.
(ii) If $\mu$ has bidegree $(q,q)$, then the current $w$ in (vii) has bidegree $(k+\ell+q-1,k+\ell+q-1)$. Since $w$ is the pushforward of a current on $Z$, it vanishes if $k+\ell+q-1<{\rm codim\,}Z$. ∎
@@
Now (ii) follows since it holds when the metric is smooth, and hence also when it is qas and non-degenerate by Theorem 6.5 (ii).
Notice that a holomorphic map $\tau\colon X^{\prime}\to X$ induces a holomorphic map $\tau\colon{\mathbb{P}}(\tau^{*}E)\to{\mathbb{P}}(E)$. Part (iii) now follows from Definition 6.1 and Proposition 5.2 (iii).
@@
If the metric on $E$ is identically 0, i.e., $Z^{\prime}={\mathbb{P}}(E)$, then ${\bf s}(E,h,h_{0})={\bf s}(E,h_{0})$ and then the statement is trivial. Let us thus assume that $Z^{\prime}$ is not equal to ${\mathbb{P}}(E)$. It may very well happen anyhow that $Z=X$. Notice that $p^{*}\gamma$ is a product as in (4.1) on ${\mathbb{P}}(E)$. Let $\psi$ and $\psi_{0}$ be as above. Then $\psi-\psi_{0}$ is a global function on ${\mathbb{P}}(E)$ and, cf. Proposition 4.8,
@@
We now consider (v). The metrics $h^{\prime}$ and $h_{\epsilon}=h+{\epsilon}h_{0}$ are both qas and non-degenerate. By Theorem 6.5 (iii) therefore
(6.28) ${\bf s}(E,h+{\epsilon}h_{0})\wedge{\bf s}(E^{\prime},(1+{\epsilon})h^{\prime})\wedge\mu={\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h+{\epsilon}(h^{\prime}\oplus h_{0}))\wedge\mu.$
@@
Since $s(E^{\prime},(1+{\epsilon})h^{\prime})=s(E^{\prime},h^{\prime})$, (i) now implies that the left hand side of (6.28) tends to ${\bf s}(E,h)\wedge{\bf s}(E^{\prime},h^{\prime})\wedge\mu$ and the right hand side tends to ${\bf s}(E^{\prime}\oplus E,h^{\prime}\oplus h)\wedge\mu$.